Find the probability of drawing a diamond card in each of the consecutive draws from a well shuffled pack of cards, if the card drawn is not replaced after the first draw.
Answer by ewatrrr(24785)

P(drawing a diamond card in each of the consecutive draws )
card drawn is not replaced
P = (13/52)(12/51)(11/50)(10/49)...(4/43)(3/42)(2/41)(1/40)
........
Re TY: IF you are wishing to find the Probability of drawing ALL 13 Diamonds
consecutively: Yes, P would be basically = Zero of doing that.
P(drawing 13 Diamonds consecutively) = 13!/( 52!/39!) = 1.57476952244911E-12
.........
generally: Questions concern a 5-card Flush, for ex.
P = (13/52)(12/51)(11/50)(10/49)(9/48) = 4.952x10^(-4) rounded